I'm a little confused on this question.
I don't know how to type a square root, but inside the square root is t+30=t
How does this problem get worked?
Thanks

Answer by mananth(16946)

sqrt(t+30)=t
square both sides
t+30=t^2
t^2-t-30=0
t^2-6t+5t-30=0
t(t-6)+5(t-6)=0
(t-6)(t+5)=0
t=6 OR t=-5